OpenStudy (anonymous):

complete the series: 3760, 1880, 360, 180, 45, ? A. 13 B. 15 C. 11.5 D. 22.5 plz explain

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@dan815 @paki @Ashleyisakitty

OpenStudy (anonymous):

is it D?

OpenStudy (tkhunny):

Hate, hate, hate, HATE these problems. There are infinitely many answers. 3760, 1880, 360, 180, 45, ? Try dividing successive pairs... 3760/1880 = 2 1880/360 = 5 2/9 360/180 = 2 180/45 = 4 That doesn't seem to be getting us anywhere. Try subtracting successive values... 3760 - 1880 = 1880 1880 - 360 = 1520 360 - 180 = 180 180 - 45 = 135 That doesn't look very helpful, either. Just play around until you find something,

OpenStudy (anonymous):

ahhhhhhhhhhhh

OpenStudy (anonymous):

some say that it's D coz 45/2=22.5 but that doesnt make sense

OpenStudy (tkhunny):

Not every sequence males a ton of sense. You just have to find something. Seriously, though, you can put ANY number next and it can be explained.

OpenStudy (anonymous):

lol it's from a university entry test so u can't put any no. coz every que matters in that and u can't risk losing a mark

ganeshie8 (ganeshie8):

read it like this : 3760 --> 1880 360 ---> 180 45 ----> ?

OpenStudy (anonymous):

then it'll be 22.5...that makes sense @ganeshie8

OpenStudy (anonymous):

thanx

OpenStudy (tkhunny):

That's one way to look at it - simply ignore any connection between 1880 and 360. This might be described as an overlapping pair of sequences.

ganeshie8 (ganeshie8):

the first terms are coming from nowhere, but the second terms are generated by dividiing... again there are plenty of mistakes in this post. few are below : 1) its a sequence, not series 2) you cant describe a sequence just by listing first few terms

OpenStudy (anonymous):

its an iq que from last year past paper for entry test...i was preparing and couldn't figure it out so i posted it here as it is @ganeshie8

ganeshie8 (ganeshie8):

its okay, the question as it stands is not mathematically meaningful. but yeah your brain somehow manages to see the intent of the author of question

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

@ganeshie8 care for confirming my ans for some other iq ques?

ganeshie8 (ganeshie8):

il try.. ask

OpenStudy (tkhunny):

"your brain somehow manages to see the intent of the author of question" No, that just isn't the case. Maybe, with some great stroke of luck, you might manage to trip over the intent of the author. You can't know it. Even if you find it, you can't claim any directed path to the solution.

OpenStudy (anonymous):

1.d or c (most probably d) 2.a 3.a 4.b 5.c

OpenStudy (anonymous):

tell me right one with reason if i did any wrong? @ganeshie8 @tkhunny

ganeshie8 (ganeshie8):

what are your reasons for picking them ?

OpenStudy (anonymous):

umm @ganeshie8 i dont have much time to tell the reasons

OpenStudy (anonymous):

@ganeshie8 @tkhunny check these too 6.a 7.d 8.d

ganeshie8 (ganeshie8):

what do u mean u dont have much time, are u giving a test ?

OpenStudy (tkhunny):

Are you actually in the exam as we speak? This would be cheating.

OpenStudy (anonymous):

no i got university entry test + my college exams in december and i got so much to prepare for thats y i don't have a min to waste

OpenStudy (anonymous):

the ques i just solved r from this site http://entrytest-preparation.blogspot.com/2014/02/questions-of-iq-from-net-1-2014.html

OpenStudy (anonymous):

@ganeshie8 @tkhunny

OpenStudy (anonymous):

looks like ur not gonna check my ans

ganeshie8 (ganeshie8):

#7 : d because you can't find time in calendar #8 : d (each column adds up to 90) #9 : d #10 : a (multiply top and bottom by 5)

OpenStudy (anonymous):

each colum adds upto 99 :P

OpenStudy (anonymous):

what bout 1,2,3,4,5,6?

ganeshie8 (ganeshie8):

#1 : d (all other branches study non living stuff)

ganeshie8 (ganeshie8):

#2 : a (all others are more related to yeast)

OpenStudy (anonymous):

better reason for 2 is that the rest r baked products

ganeshie8 (ganeshie8):

probably as biscuits are not so yeasty in many parts of the world

OpenStudy (anonymous):

thats i i said 'baked'